OpenStudy (anonymous):

At each point (x,y) on the curve the slope is 2x+1; the curve passes through the point (-3,0). Find the equation of the curve.

OpenStudy (anonymous):

y=x + x^2 -6 y'=2x+1 y(0) 0=-3 +9 -6

OpenStudy (anonymous):

There you go...! I forgot that the derivative gives you the slope of the eqn that you have. Thanks!